OpenStudy (o-b-server):

Please answer soon! Which are solutions to y = x - 4? (6,3) (4,0) (3, -1) (2,-4) Choose all correct answers

OpenStudy (mirandaregina):

First, you rearrange the equation so it looks like this: 4 - x = y. From there, you take one of the points you have listed and plug them into the equation. I'll do (4, 0) for example: 4 - 4 = 0. This is a true statement, so it is a solution to y = x - 4.

OpenStudy (o-b-server):

Ah okay thank you very much!